# Astrophysical signals from the neighboring brane

### v2

@@ 3.2 The Planck mass
[…]
where $v_{n-1}$ is the volume of an $n-1$-dim sphere of unit radius
$v_{n-1}=\frac{\pi^{\frac{n-1}{2}}}{\Gamma(\frac{n-1}{2}+1)}.$ (13)
The Planck mass $m_{4}$ is measured at the low energies, and we assume that the inequality $R_{4}\ll R_{2}$ holds.
@@
The volume (13) has maximum $v_{n-1}\simeq 2500$ at $n\simeq 20$ which gives the estimation of the smallest value of $m_{D}\sim 0.02m_{4}\simeq 10^{17}$eV. Here and below we suppose $m_{D}\sim 10^{17}\div 10^{18}$eV and the scale of the two-brane structure $l_{extra}\sim 10^{-14}\div 10^{-15}$eV-1 for the estimations. Other terms in (3.2) contribute to the vacuum energy density $\Lambda$, see discussion in [Bronnikov:2023lej] , which is not used here. The knowledge of the metric functions $\gamma(u),r(u)$ as in Figure 1 permits to restore the 4D Planck mass $m_{D}$, see expression (14).
A non-trivial scenario arises if the number $n$ of extra dimensions is large. In this case, the volume of extra dimensions (13) is a rapidly decreasing function of the number of extra dimensions. For example, $v_{n}\simeq 4\cdot 10^{-15}$ at $n=100$, so
$m_{D}\sim 10^{8}m_{4}\sim 10^{27}\text{GeV}$
The onset of the quantum regime occurs at a scale $1/m_{D}\sim 10^{-8}/m_{4}$, which is significantly smaller than the typically mentioned estimation $1/m_{4}$.
@@ 4 Fields splitting between two branes
where $\{Y_{k}(y)\}$ is a set of orthonormalized functions. It is usually assumed that they satisfy the equation
@@
4.1 Fermions distribution over the extra dimensions. Analytical solution
In this subsection we shortly remind the results of [Popov:2024nax] concerning the distribution of fermions along the radial extra-dimensional coordinate $u$ near the brane and in the bulk. General form of the free massless fermion action is
@@
The mass term is assumed to be generated by the interaction with the Higgs field. The particular form of the curved gamma matrices $\Gamma^{M}$ and the covariant derivatives can be found in [Oda_2000]. The useful discussion is also given in [Appelquist:2000nn, Randjbar-Daemi:2000lem]. The more thorough model in even dimensions is outlined in [Wan:2023usr]. That paper studies conditions for decoupling the left and right chiralities of spinors. A useful discussion is also given in [Appelquist:2000nn, Randjbar-Daemi:2000lem]. A 6D spinor
$\Psi=\left(\begin{array}[]{c}\xi_{1}\\ \xi_{2}\end{array}\right)$ (18)
acting in six dimensions is equivalent to a pair of 4-dimensional Dirac spinors $\xi_{1}$ and $\xi_{2}$. As shown in [Gogberashvili_2007], both 4D spinors, $\xi_{1}$ and $\xi_{2}$, are equal in the s-wave approximation used in our approach. In this paper, we are interested in the spinor distribution over the extra dimensions. The spinor $\psi$ is characterized by both spinor and group indices and functions within a six-dimensional framework. The field decomposition (15) in this case reads
$\Psi(x,y)=\psi(x)Y_{f}(u).$ (19)
Here the normalized function $\psi(x)$ is endowed with spinor and group indices acting in 4-dimensional space.
This study focuses on the distribution of fields across the extra-dimensional coordinates rather than their reduction to four-dimensional physics. Consequently, our attention is directed towards the form of the function $Y_{f}(u)$ which is a solution to equation
@@
The analytical solution to this equation has the form
$Y_{f}(u)=c_{2}r(u)^{-1/2}e^{-2\gamma(u)}.$ (21)
@@
The particular distribution over the radial coordinate $u$ is shown in Fig.2. The more detailed analysis is presented in [Popov:2024nax]. The distribution has the sandwich-type structure with sharp peaks at the coordinates $u_{1}$ and $u_{2}$ and almost zero field value between them. The region proximate to the coordinate $u_{1}$ is designated brane-1, and the area near the coordinate $u_{2}$ is labeled brane-2.
4.2 Spreading a field over two branes and splitting it into two independent effective fields.
@@
Here $\lambda_{in}$ is an initial set of physical parameters. An action $S$ in (22) is a functional of a function $\zeta$ acting in D-dim. At the highest energy levels, where $H\sim m_{D}$, this field is dispersed throughout the entire extra space. However, as the system’s energy decreases to lower levels, the scenario changes. As was discussed in the subsection 2.1, the extra space structure is formed at the scale $l_{e}\lesssim 10^{-27}$cm. At the present time, the smallest scale where achieved at the colliders $l\sim 10^{18}$cm. The inequality $l_{e}\ll l_{s}$ allows to neglect the 4D variation of fields and use the classical equations for fixing the extra-dimensional metric and the fields distribution over extra coordinates $y$. The decomposition (15) in the form
@@
where the main order of the approximation is used in the last line. Knowledge of the function $Y_{cl}(y)$ permits to integrate over the extra dimensional coordinate $y$ which lead to the renormalization of the initial physical parameters values $\lambda_{in}\to\lambda_{eff}$. $S^{(4)}$ in effective 4D action. The general conclusion is that the low energy effective parameters $\lambda_{eff}$ appears to be dependent on accidental field distribution at high energies. This result being very important, has been performed for various fields in [Rubin:2016ude, Petriakova:2023klf, Bronnikov:2023lej, Popov:2024nax].
@@
The first step in application of the Method of Stationary Phase is to find classical function delivering an extrema of action. Assume that this function has two sharp maxima $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ and is (almost) zero between them region $\mathcal{U}_{0}$, like in Fig.2. Then, the action can be segmented into three components based on integration over the regions $\mathcal{U}_{1}$, $\mathcal{U}_{0}$, and $\mathcal{U}_{2}$ Taking all this into consideration, the transition amplitude (22)
@@
The fields in the different branes do not feel each other. However, there are waves that could transfer energy between the branes. At the low energies, the wavelengths of external excitations are much larger compared to the extra space scale and hence cannot be excited.
4.3 The effective parameters of a field are brane-dependent.
@@
The same is true for fields concentrated near branes. The field near one brane does not ’’feel’’ the field located on the other brane due to the smallness of the overlapping integral. This is true for the zero mode $y_{cl}$. The KK mode excitations are forbidden if their typical wavelength $l\gg l_{extra}$,  ($l_{extra}$ is the distance between the branes) is much smaller than the electroweak scale.
@@
4.4 Electromagnetic Field and Gravity between branes
In this subsection, we use the results obtained in [Cui:2020fiz, Popov:2024nax] to show that the electromagnetic field and the gravitational excitations are distributed uniformly between the branes.
We start with the action of the abelian gauge field and the standard form of its D-dimensional action, see e.g. [Dubovsky:2000av]
$S_{gauge}=-\frac{\text{g}_{D}}{4}\int d^{D}X\sqrt{|g_{\text{D}}|}g^{PM}g^{QN}F_{PQ}F_{MN},$
$F_{QN}=\partial_{Q}A_{N}-\partial_{P}A_{Q}.$ (34)
Here $\text{g}_{D}$ is the D-dimensional coupling constant. The classical equations
$\partial_{P}(\sqrt{|g_{D}|}g^{PQ}g^{MN}F_{QN})=0$ (35)
are used to find the field distribution along the extra coordinate $u$. Keeping in mind expression (15) in the form
$A^{N}(x,u)=a^{N}(x)Y_{em}(u)$ (36)
equations (35) can be rewritten as follows ($n=2$)
$M=\mu:\quad a^{\mu}\left[Y_{em}^{\prime\prime}+\left(2\gamma^{\prime}+\frac{r^{\prime}}{r}\right)Y_{em}^{\prime}\right]=0,$
$M=\theta:\quad a^{\theta}\left[Y_{em}^{\prime\prime}+\left(4\gamma^{\prime}-\frac{r^{\prime}}{r}\right)Y_{em}^{\prime}\right]=0.$ (37)
Here, only nontrivial equations are written. Remind that the conditions $\partial_{\mu}\cdot=0,\,\mu=0,1,2,3$ are used. The latter holds at the low energies where 4D wavelengths are much larger than the extra space scale. Evidently, $Y_{em}(u)=C_{em}=const$ is the unique solution to system (4.4) if $a^{\mu},\,a^{\theta}\neq 0$ which means that the electromagnetic field is uniformly distributed between the branes. Charges located on one brane can emit the photons which are absorbed by charges on the other brane, see discussion in section 5.
Let us now discuss the gravitational perturbation over the extra coordinate $u$. The metric containing the tensor gravitational perturbations is of the form
$ds^{2}=e^{2\gamma(u)}(\eta_{\mu\nu}+h_{\mu\nu})dx^{\mu}dx^{\nu}-du^{2}-r(u)^{2}d\phi^{2}$ (38)
$\eta_{\mu\nu}=diag(1,-1,-1,-1),\quad h_{\mu\nu}\ll\eta_{\mu\nu}$ where the perturbation $h_{\mu\nu}$ represents the tensor mode obeying the transverse-traceless gauge conditions
$\eta_{\mu\nu}h^{\mu\nu}=0,\quad\partial_{\mu}h_{\mu}^{\nu}=0.$
During this paper, we are interested in the main term of the decomposition (15) and represent the perturbations in the form
$h_{\mu\nu}(x,u)=H^{(0)}_{\mu\nu}(x)Y_{0}(u)$ (39)
Our aim is to find the distributions over the extra dimensions so that we put $H^{(0)}_{\mu\nu}(x)=const$ (it is assumed that the $x-$derivatives are small, or $\partial_{\mu}\cdot=0$ ). After some algebra, see [Cui:2020fiz], equations for the linear tensor perturbations are reduced to a non-trivial equation
$\left[\left(4\gamma^{\prime}+\frac{r^{\prime}}{r}\right)\partial_{u}+\partial_{u}^{2}+\frac{\partial_{u}f_{R}}{f_{R}}\partial_{u}\right]Y_{0}(u)=0.$ (40)
for the tensor mode (39). One of the possible solutions of the equation (40) is $Y_{0}=const$, as it should be for the zero mode. This means that the gravitational field is uniformly distributed between the two branes. Matter on one brane experiences the gravitational forces produced by matter on the other brane.
5 Manifestation of the second brane
In this section, we will briefly examine the possible observational implications of the hypothetical sandwich model of extra dimensions discussed in this paper. In the previous section, we have shown that
- both branes are filled with similar types of fields;
- the photons and the gravitational waves in 4D are distributed uniformly over the extra dimensions;
- The hierarchy problem might be solved exclusively within a single brane framework, implying that the formation of complex astrophysical structures — such as stars — could be restricted to brane-1.
@@ 5.1 Photon interaction
The subsection 2.2 deals with the challenges of interpreting these particles as dark matter candidates. The most obvious property of dark matter is its extremely weak interaction with the electromagnetic field. Here we will answer two questions: are the charges of the electrons on different branes equal, and can such electrons form a bound state? So, we have to consider a D-dim extension of the SM action for the electron-photon interaction
[…]
Here $\alpha_{D}$ is a coupling constant in D dimensions and $A_{N}$ is the gauge field. It is shown above that the gauge field is uniformly distributed over the extra dimensions, i.e. $A_{N}(x,y)=C_{em}a_{N}(x)$. Action (41) can be reduced to the 4-dim action:
@@
where expressions (19) and (32) have been kept in mind. We also used the fact that the extra space distribution of the fermions is included into the normalization of the wave functions $\psi_{1,2}$ and $\alpha=C^{2}_{em}\alpha_{D}$. The observable coupling constant $\alpha$ is the same on both branes. The parameter $\alpha_{12}$ is proportional to the overlapping integral (33) and hence, is extremely small.
@@
The third term represents the vertex containing both types of electrons in conjunction with the electromagnetic field. The transition of the heavy electron to another brane, accompanied by photon emission, is associated with this vertex. The latter could also lead to the bound states containing the two electrons on different branes. Both of these effects are negligibly small due to the smallness of the coupling constant $\alpha_{12}$.
5.2 The Higgs mechanism on different branes
@@
$-\left[\partial^{2}_{u}+\left(4\gamma^{\prime}+\dfrac{r^{\prime}}{r}\right)\partial_{u}\right]Y_{H}(u)=\nu Y_{H}(u)-\lambda\,Y_{H}^{3}(u),$ (45)
As was shown in [Bronnikov:2023lej], the field distribution over the extra dimensions is specific to each causally separated 3D region below the horizon. Consequently, the field distributions are unique for each universe generated during the de Sitter stage. One of the solutions is shown in Fig. 3.
Figure 3: The Higgs field distribution over the internal coordinate $u$ is specific to each brane. The background metric functions are represented in Fig.1. Additional conditions are $Y(1.5)=0,\,Y^{\prime}(1.5)=10^{-2}$.
The expressions for the Higgs mass and the coupling constant as functionals of the function $Y_{H}(u)$ were obtained in [Bronnikov:2023lej].
[…]
Here $\nu,\,\Lambda$ are the initial physical parameters of the Higgs field in D-dim, ${v}_{\text{n}-1}$ is the volume of the $(\text{n}-1)$-dimensional sphere. This is the specific example of the renormalization of the initial physical parameter values $\lambda_{in}\to\lambda_{eff}$ introduced in subsection 4.2, where $\lambda_{in}=\nu,\,\Lambda$ and $\lambda_{eff}=m_{h},\,\lambda_{h}$.
The Higgs distribution $H$ can be split as $H(x,y)=H_{1}(x,u)+H_{2}(x,u)$, see general expression (32). Recall that the distributions $H_{1}(x,u)$ and $H_{2}(x,u)$ act as independent fields as is discussed in 4.2 and 4.3. The function $H_{1}(x,u)$ decreases rapidly far from the endpoint $u_{1}$ so that the observer on the brane-1 perceives the Higgs field as $H(x,y)=H_{1}(x,u)=h_{1}(x)Y^{(1)}_{H}(u)$. The function $H_{2}(x,u)$ quickly decreases far from the endpoint $u_{2}$ and a possible observer on the brane-2 perceives the Higgs field as $H(x,y)=H_{2}(x,y)=h_{2}(x)Y^{(2)}_{H}(u)$. Therefore, the 4D parameters of the Higgs field $m_{h},\lambda_{h}$ are functionals of the distribution $Y^{(1)}_{H}(y)$ for the brane-1 observer and $Y^{(2)}_{H}(y)$ for the brane-2 possible observer. Hence, they are different on both branes. The same is true for the Higgs vacuum average
@@
It was declared from the beginning that the model is free from initially small physical parameter values. This means that the integrals as in (46) and (47) are of the order of unity if no special effort was made. The fine-tuning should be done in the expression (46) to obtain the observable Higgs mass and its vacuum average. As shown in the paper [Bronnikov:2023lej], the vacuum average can be adjusted to the known value $v_{h}\sim 10^{-17}m_{Planck}$ by an appropriate selection of the classical distribution $Y_{H}(u)$ over the extra dimensions. In the framework of the two-brane model discussed here, this procedure can be done for the brane-1 where the observer is assumed. This means that $Y_{H}(u)$ should be substituted by $Y^{(1)}_{H}(u)$ in formulas (46) and (47). The Higgs vacuum average on the brane-2 remains uncertain.
@@
of the fermions in D-dim on brane-1 (marked as $Y^{(1)}_{L,R}$) and brane-2 (marked as $Y^{(2)}_{L,R}$) leads to the following action
@@
If we live on brane-1, the SM in 4 dimensions and the parameter $f^{(1)}_{e}$ must have the observable value $f^{(1)}_{e}=m_{e}/v^{(1)}_{h}\sim 10^{-6}$, see [Workman:2022ynf]. This value is very small and hence needs to be fine-tuned in addition to the vacuum average $v_{h}$. This can be done by selecting appropriate functions $Y^{(1)}_{L,R}(u)$ in (53).

### v3

@@ 3.2 The Planck mass
$m_{4}^{2}=m_{D}^{D-2}v_{n-1}\int due^{2\gamma(u)}r(u)f_{R}(R_{2}+\mathcal{R}),$ (11)
where $v_{n-1}$ is the volume of an n1n-1-dim sphere of unit radius
$v_{n-1}=\frac{2\pi^{\frac{n}{2}}}{\Gamma(\frac{n}{2})}.$ (12)
The Planck mass $m_{4}(H=0)=m_{Pl}$ is measured at the low energies, where we assume that the inequality $R_{4}\ll R_{2}$ holds.
@@
The volume (12) has maximum $v_{n-1}\simeq 33$ at $n\simeq 7$ which gives the estimation of the smallest value of $m_{D}\sim 0.2m_{4}\simeq 10^{18}$GeV. Here and below we suppose $m_{D}\sim 10^{17}\div 10^{18}$GeV and the scale of the two-brane structure $l_{extra}\sim 10^{-14}\div 10^{-15}$GeV-1 for the estimations. Other terms in (3.2) contribute to the vacuum energy density $\Lambda$, see discussion in [24] , which is not used here. The knowledge of the metric functions $\gamma(u),r(u)$ as in Figure 1 permits to restore the 4D Planck mass $m_{D}$, see expression (11).
A non-trivial scenario arises if the number $n$ of extra dimensions is large. In this case, the volume of extra dimensions (12) is a rapidly decreasing function of the number of extra dimensions. For example, $v_{n}\simeq 6.5\cdot 10^{-4}$ at $n=3$, so
$m_{D}\sim 10^{2}m_{4}\sim 10^{21}\text{GeV}.$
The onset of the quantum regime occurs at a scale $1/m_{D}\sim 10^{-2}/m_{4}$, which is significantly smaller than the typically mentioned estimation $1/m_{4}$.
@@ 4 Fields splitting between two branes
where $\{Y_{k}(y)\}$ is a set of orthonormalized functions describing the extra metric excitations which are neglected if the energy is not extremely high. The normalization constant is hidden in the field $\Upsilon(x)$ and is used at the final state to convert the 4D kinetic term to its standard form, for example $\frac{1}{2}\partial_{\mu}\Upsilon(x)\partial^{\mu}\Upsilon(x)$, in case of the scalar field. It is usually assumed that they satisfy the equation
@@
4.1 Fermions distribution over the extra dimensions.
In this subsection we shortly remind the results of [61] concerning the distribution of fermions along the radial extra-dimensional coordinate $u$ near the brane and in the bulk. General form of the free massless fermion action is
@@
The mass term is assumed to be generated by the interaction with the Higgs field. The particular form of the curved gamma matrices $\Gamma^{M}$ and the covariant derivatives can be found in [58]. The useful discussion is also given in [73, 5, 64].
As shown in [42], a 6D spinor in the s-wave approximation
$\Psi=\left(\begin{array}[]{c}\xi_{1}(x^{\mu})\\ \xi_{2}(x^{\mu})\end{array}\right)Y_{f}(u)\equiv\psi(x)Y_{f}(u)$ (17)
is equivalent to a pair of 4D Dirac spinors, $\xi_{1}$ and $\xi_{2}$, which are equal in this case. Here the normalized function $\psi(x)$ is endowed with spinor and group indices acting in 4D space. This study focuses on the distribution of fields $Y_{f}(u)$ across the extra-dimensional coordinates rather than their reduction to four-dimensional physics. Consequently, our attention is directed towards the function $Y_{f}(u)$ in the form
$Y_{f}(u)=r(u)^{-1/2}e^{-2\gamma(u)},$ (18)
see Fig.2, which is a solution to equation
@@
This equation was thoroughly derived in [58].
@@
The particular distribution over the radial coordinate $u$ is shown in Fig.2. This distribution has two sharp peaks at the coordinates $u_{1}$ and $u_{2}$ as can be seen also from the analytical expression (18) and Fig. 1. The region proximate to $u_{1}$ is designated as brane 1, and the area near $u_{2}$ is labeled as brane 2. The fields on both branes act independently because the distribution between them is almost zero. A rigorous proof of this statement is in the next subsection. The field behavior on brane 1 and brane 2 is described by the distribution in Fig. 2, spanning the intervals $(0,\sim 1)$ and $(\sim 2,\pi)$, respectively.
An important fact is that the observable fermions are described by the wave function $\psi(x)$ in the decomposition $\Psi(x,u)=\psi(x)Y_{f}(u)$ and propagate in the four-dimensional spacetime orthogonal to the extra coordinate $u$. The function $Y_{f}(u)$ exhibits two distinct peaks near the endpoints $u_{1,2}$. Thus, the fermion field density is proportional to $|Y_{f}(u)|^{2}$ and is maximal near the endpoints $u_{1,2}$.
Also, $Y_{f}(u)\to 0$ (or is approximately vanishing) in the interbrane region, so that the fermion field density tends to zero there. As a result, the fermion excitations $\psi(x)$ on each brane do not influence each other and can be treated as independent. A rigorous proof follows in the next subsection.
4.2 Fields distribution in branes
@@
Here $\lambda_{in}$ is an initial set of physical parameters. An action $S$ in (20) is a functional of a function $\zeta$ acting in D-dim. At the highest energy levels, where $H\sim m_{D}$, this field is dispersed throughout the entire extra space. However, as the system’s energy decreases to lower levels, the scenario changes. As was discussed in the section 2, the extra space structure is formed at the scale $l_{e}\lesssim 10^{-27}$cm. At the present time, the smallest scale where achieved at the colliders $l\sim 10^{18}$cm. The inequality $l_{e}\ll l_{s}$ allows to neglect the 4D variation of fields and use the classical equations for fixing the extra-dimensional metric and the fields distribution over extra coordinates $y$. The decomposition (14) in the form
@@
where the main order of the approximation is used in the last line. Knowledge of the function $Y_{cl}(y)$ permits to integrate over the extra dimensional coordinate $y$ which lead to the renormalization of the initial physical parameters values $\lambda_{in}\to\lambda_{eff}$ in effective 4D action $S^{(4)}$. The general conclusion is that the low energy effective parameters $\lambda_{eff}$ appears to be dependent on accidental field distribution at high energies. This result being very important, has been performed for various fields in [67, 60, 24, 61].
@@
4.3 Splitting a field into two independent effective fields
Here, we demonstrate that a field distributed across both branes can effectively be treated as two independent fields. The first step in application of the Method of Stationary Phase is to find classical function delivering an extrema of action. Assume that this function has two sharp maxima $\mathcal{U}_{1}$ and $\mathcal{U}_{2}$ and is (almost) zero between them region $\mathcal{U}_{0}$, like in Fig.2. Then, the action can be segmented into three components based on integration over the regions $\mathcal{U}_{1}$, $\mathcal{U}_{0}$, and $\mathcal{U}_{2}$ Taking all this into consideration, the transition amplitude (20)
@@ 4.3 Splitting a field into two independent effective fields
the two localized fields interact negligibly due to exponentially suppressed overlap. However, there are waves that could transfer energy between the branes. At the low energies, the wavelengths of external excitations are much larger compared to the extra space scale and hence cannot be excited.
4.4 Weak interaction of particles localized in different branes
@@
The same is true for fields concentrated near branes. The field near one brane interacts extremely weakly with the field in the other brane due to the smallness of the overlap integral. This is true for the zero mode $y_{cl}$. The KK mode excitations are forbidden if their typical wavelength $l\gg l_{extra}$,  ($l_{extra}$ is the distance between the branes) is much smaller than the electroweak scale.
@@ 4.5 Photon interaction
In this subsection, we use the results obtained in [32, 61] where it was shown that the electromagnetic field and the gravitational excitations are distributed uniformly between the branes.
Here we address two questions: (1) Is the electric charge of electrons identical across different branes? (2) Can electrons transit between branes? So, we have to consider a D-dim extension of the SM action for the electron-photon interaction
$S_{int}\equiv\alpha_{D}{v}_{\text{n}-1}\int\sqrt{|g_{4}|}d^{4}x\int_{u_{1}}^{u_{2}}du\sqrt{|g_{n}|}\bar{\Psi}\Gamma^{N}A_{N}\Psi.$ (32)
Here $\alpha_{D}$ is a coupling constant in D dimensions and $A_{N}$ is the gauge field. It is shown in [61] that the gauge field is uniformly distributed over the extra dimensions, i.e. $A_{N}(x,y)=Const\cdot a_{N}(x)$. Action (32) can be reduced to the 4-dim action:
@@
where expression (30) have been kept in mind. We also used the fact that the extra space distribution of the fermions is included into the normalization of the wave functions $\psi_{1,2}$ and $\alpha=Const^{2}\alpha_{D}$. The observable coupling constant $\alpha$ is the same on both branes. The parameter $\alpha_{12}$ is proportional to the overlapping integral (31) and hence, is extremely small.
@@
The third term represents the vertex containing both types of electrons in conjunction with the electromagnetic field. Hence, the transition of the heavy electron to another brane, accompanied by photon emission, is associated with this vertex. This effect is negligibly small due to the smallness of the coupling constant $\alpha_{12}$.
4.6 The Higgs mechanism in different branes
@@
$\left[\partial^{2}_{u}+\left(4\gamma^{\prime}+\dfrac{r^{\prime}}{r}\right)\partial_{u}\right]Y_{H}(u)=\nu Y_{H}(u)-\lambda\,Y_{H}^{3}(u),$ (36)
as was shown in [24]. One of the solutions shown in Fig. 3 indicates that the Higgs field also exhibits sharp maxima on the branes.
Figure 3: The Higgs field distribution over the internal coordinate $u$ is specific to each brane. The background metric functions are represented in Fig.1. Additional conditions are $Y_{H}(1.5)=0,\,Y_{H}^{\prime}(1.5)=10^{-2}$. The sharp peaks which define the branes are the result of numerical solutions to equation (36) in wide range of parameters. The Higgs field is approximately zero in the interval between two black points. The fermion fields are also concentrated near the branes, see Fig.2.
The Higgs distribution $H$ can be split as $H(x,y)=H_{1}(x,u)+H_{2}(x,u)$, see general expression (30). Recall that the distributions $H_{1}(x,u)$ and $H_{2}(x,u)$ act as independent fields as is discussed in 4.2 and 4.4. The function $H_{1}(x,u)$ decreases rapidly far from the endpoint $u_{1}$ so that the observer on the brane-1 perceives the Higgs field as $H(x,y)=H_{1}(x,u)=h_{1}(x)Y^{(1)}_{H}(u)$, see Fig.3 left part of the plot. The function $H_{2}(x,u)$ quickly decreases far from the endpoint $u_{2}$ and a possible observer on the brane-2 perceives the Higgs field as $H(x,y)=H_{2}(x,u)=h_{2}(x)Y^{(2)}_{H}(u)$.
The expressions for the Higgs mass and the coupling constant as functionals of the function $Y_{H}(u)$ were obtained in [24]. Applied to the brane-1, these parameter are as follows
$m^{2}_{h}={v}_{\text{n}-1}\int_{u_{1}}^{u_{1}+\Delta}\Bigl(-(\partial_{u}Y_{H})^{2}+\nu\,Y_{H}^{2}(u)\Bigr)\text{e}^{4\gamma(u)}r^{\text{n}-1}(u)\,du,$ (37)
$\lambda_{h}={v}_{\text{n}-1}\int_{u_{1}}^{u_{1}+\Delta}\lambda\,Y_{H}^{4}(u)\,\text{e}^{4\gamma(u)}r^{\text{n}-1}(u)\,du,\quad{v}_{\text{n}-1}\equiv\int d^{n-1}y\sqrt{|{g}_{n-1}|}=\dfrac{2\pi^{\tfrac{\text{n}}{2}}}{\,\Gamma\left(\tfrac{\text{n}}{2}\right)}.$ (38)
Here $\nu,\,\Lambda$ are the initial physical parameters of the Higgs field in D-dim, ${v}_{\text{n}-1}$ is the volume of the $(\text{n}-1)$-dimensional sphere.
The effective width of the brane, denoted by $\Delta$, is defined as the characteristic size of the extra dimension where the energy density of the localized fermion and the Higgs fields are non-negligible. The interaction between the Higgs field $H$ and the fermions $\Psi$ is described by vertex (40). Since the Higgs field and the fermion field $\Psi$ are tightly localized near the brane $u_{1}$ (e.g., $u_{1}=0$), the main contribution to the integral over the extra dimension is restricted to the interval $(u_{1},u_{1}+\Delta)$. Therefore, the 4D parameters of the Higgs field $m_{h},\lambda_{h}$ are functionals of the distribution $Y^{(1)}_{H}(u)$ for the brane-1 observer, see Fig.3. The same is true for the Higgs vacuum average
@@
These arguments can be applied to the brane-2 with the substitutions $Y^{(1)}_{H}(u)\to Y^{(2)}_{H}(u)$ and the interval of integration $(u_{1},u_{1}+\Delta)\to(u_{2}-\Delta\to u_{2})$ in expressions (37) and (38). Hence, the vacuum average is different on both branes.
It was declared from the beginning that the model is free from initially small physical parameter values. This means that the integrals as in (37) and (38) are of the order of unity if no special effort was made. The fine-tuning should be done in the expression (37) to obtain the observable Higgs mass and its vacuum average. As shown in the paper [24], the vacuum average can be adjusted to the known value $v_{h}\sim 10^{-17}m_{Planck}$ by an appropriate selection of the classical distribution $Y_{H}(u)$ over the extra dimensions. In the framework of the two-brane model discussed here, this procedure can be done for the brane-1 where the observer is assumed. This means that $Y_{H}(u)$ should be substituted by $Y^{(1)}_{H}(u)$ in formulas (37) and (38). The Higgs vacuum average on the brane-2 remains uncertain.
@@
[…]
@@
If we live on brane-1, the SM in 4 dimensions and the parameter $f^{(1)}_{e}$ must have the observable value $f^{(1)}_{e}=m_{e}/v^{(1)}_{h}\sim 10^{-6}$, see [74]. This value is very small and hence needs to be fine-tuned in addition to the vacuum average $v_{h}$. This can be done by selecting appropriate functions $Y^{(1)}_{L,R}(u)$ in (44).
@@
4.7 Stability
The study of static solutions stability in extra dimensional models based on nonlinear $f(R)$ gravity is extremely complicated problem. The results obtained up to now are based on specific assumption on the analytical form of solutions.
Tensor-mode stability for the brane background is analyzed in [76, 75, 54]. The stability of the compact static metric is usually connected to the radion field [52] which is defined as the physical distance between two branes which leading to interesting results if the metric is maximally symmetric or in the models inspired by the Rassel-Sundrum one. In the model developed here, the situation is more complicated because of the absence of appropriate metric symmetries and the numerical character of solutions interpreted as the thick branes. Our numerical simulations made in [23] indicate that perturbations are attenuated with time.
The sensitivity of the model to additional conditions is also worth noting. In this approach, the brane structure emerges from solutions to the field equations subject to specific additional conditions—namely, the metric functions and their derivatives are fixed at a particular coordinate $u$. The nontrivial fact is that vast majority of solutions terminate at certain coordinates, exhibiting the properties of static branes. A slight variation in these additional conditions leads to a correspondingly small variation in the metric functions, while the presence of the branes themselves persists. Massive, radion-like perturbations appear only at extremely high energy scales, which lie far beyond the scope of the present study.
The analogy with the four-dimensional collapse of dust to form a black hole provides strong intuitive support for our model. Indeed, the final mass of a black hole is determined by the initial distribution of dust that collapses to create it. A small variation in the initial dust distribution leads to a correspondingly small variation in the black hole mass, in direct analogy to the behavior of the extra-dimensional metric functions discussed above. Evidently, the resulting black hole masses form a continuous, parameterized set of static solutions. The same can be said of the extra-dimensional metric functions. We also note that black holes themselves represent metastable states due to Hawking radiation.
5 Matter fields in the uninhabited brane.
Problems and prospects.
In this section, we investigate potential observational signatures of our extra-dimensional model proposed earlier. As established above, the energy scale hierarchy problem represents a fundamental theoretical challenge that appears resolvable only within brane-1. This framework naturally predicts that brane-2 hosts superheavy partners to all Standard Model particles, with masses sufficiently large to: (i) serve as viable dark matter candidates, and/or (ii) potentially produce other distinctive observational signatures.
This scenario could also lead to fundamentally different astrophysical evolution in each brane: while complex structures like stars might be prohibited in brane-2, stable hydrogen-like matter could still form if QCD dynamics operate similarly across both branes.
5.1 Dark matter components?
One of the main goals of the cosmology is to understand the nature of the dark matter. Special efforts have been applied to explain the properties of this phenomenon and the origin of dark matter [27, 48, 51, 9, 14, 18]. Many theoretical models as well as the observational evidences can be found in the reviews [19, 6, 55]. The idea of extra dimensions is also attracted in the framework of this discussion, [29, 11, 49, 27]. Some of them are treated the brane idea [20, 68] which is natural basis for production of weakly interacted particles.
In this paper, we have demonstrated that brane-2 contains heavy partners of all Standard Model (SM) particles and their corresponding antiparticles. Among these, heavy neutrinos stand out as promising dark matter candidates. A detailed discussion on this topic can be found in [2, 12].
A more thorough discussion is necessary to address the implications of charged particles, such as electrons in the dark matter context. At first glance, it may seem that the dark matter cannot contain charged particles due to their significant interaction with photons (see also the discussion in [50]). This holds true for observable particles like electrons. However, the situation changes if the particles are sufficiently massive. Indeed, let us estimate the Compton effect on brane-2, i.e. the photon scattering on the non-relativistic heavy electron. This cross section depends on a particle mass $m$ as
$\sigma_{e}\sim\frac{1}{m^{2}}\sim 10^{-26}\text{cm}^{2}.$ (45)
It is easy to see that the heavy electron mass should be really large,
$m_{2}>10^{9}m_{e}\sim 10^{6}\text{GeV},$ (47)
where the value of $\sigma_{e}$ is taken from (45). The mass estimate (47) for heavy electrons on brane-2 allows one to consider them as the dark matter candidate.
Nevertheless, caution is warranted. Scattering between ordinary electrons and their massive counterparts is unsuppressed, as evidenced by the effective vertices in (34). This interaction could potentially distort the CMB and lead to unacceptable accumulation of the dark matter in the form of heavy electrons in galactic centers.
Charged particles, being extremely massive could also form bonded states like the positronium and the hydrogen. The positronium life-time $\sim m^{-1}$, ($m$ is the heavy electron mass on the brane-2) and hence decay instantly after formation. The concentration of the heavy electrons is significantly suppressed, yet it remains nonzero. Heavy hydrogen cannot be a viable dark matter candidate if strong interaction remains unchanged in brane-2.
The properties of particles in the uninhabited brane are completely unknown. While Standard Model particles exist there, their masses, coupling constants, and concentrations are undetermined. These unknown brane-2 particles form specific subdominant components of dark matter and might also explain certain observed phenomena, which could, in turn, help clarify their properties keeping in mind restrictions imposed by known and firmly established constraints from CMB photon scattering, for example. Below we perform very preliminary discussion on one of such phenomena.